a. A bond that has $1,000 par value (ace value) and a contract or coupon nderest ate of 11 percent years. The tems average tax rate is 30 percent and its marginal tax rale is 37 percent b. A new conmon stock issue that paid a $1.50 dividend last year The par value of the slock is $15, and earnings per share have grown at a rate of 8 percent per yea Thes growth rate ts expected to contnue rto the he.ฯ vatin tutiin, The company martains a constant diadendearmnas atk of 30 percent rhe price or this stock is how za, sut 6 percent halen coss ae aset ded С. internal common equity when me curent market price of the common skok is S49 The expeded didend the cong year sho ad be S3 00 incre asang here ate at an anna g se, ate of t 1 percent Te corporations tax rale s 37 percent d. A preflesred stock paying a diadend of 11 percen on a $100 par value if a new issue s offered fotation costs wil be 15 percent of the cunment price of 5178 e. A bond selling to yeid 9 percent ater totation costs, tut betore adusting for the marginal corporate tax rate of 37 percent in other words, 9 percent s the rate that equales the net proceeds from the bond with he present value of the fulure cash fows (principail and interest) A new issae would harve a floatation cost of 7 percent of the $1,135 market valuc The bonds mature in 11 a. What is the ms after tax cost of debt on the bond? 6.39 (Round to two decimai places) b. What is the cost of external common equy Round to two decimal places )
Answer #1

1: Using financial calculator

FV=1000, PV= -1135*(1-0.07)= -1055.55

PMT=11%*1000 = 110

N=11

I/Y = Pre tax cost = 10.14%

After tax cost of debt = Pre tax cost*(1-Tax)

= 10.14%*(1-0.37)

= 6.39%

2: Cost of equity = D1/Price after flotation + g

Cost of equity = 1.5*108%/ (28*(1-0.06)) + 8%

=14.16%